# Active Sensing of Knee Osteoarthritis Progression with Reinforcement Learning

### v1

@@ Abstract
Osteoarthritis (OA) is the most common musculoskeletal disease, which has no cure. Knee OA (KOA) is one of the highest causes of disability worldwide, and it costs billions of United States dollars to the global community. Prediction of KOA progression has been of high interest to the community for years, as it can advance treatment development through more efficient clinical trials and improve patient outcomes through more efficient healthcare utilization. Existing approaches for predicting KOA, however, are predominantly static, i.e. consider data from a single time point to predict progression many years into the future, and knee level, i.e. consider progression in a single joint only. Due to these and related reasons, these methods fail to deliver the level of predictive performance, which is sufficient to result in cost savings and better patient outcomes. Collecting extensive data from all patients on a regular basis could address the issue, but it is limited by the high cost at a population level. In this work, we propose to go beyond static prediction models in OA, and bring a novel Active Sensing (AS) approach, designed to dynamically follow up patients with the objective of maximizing the number of informative data acquisitions, while minimizing their total cost over a period of time. Our approach is based on Reinforcement Learning (RL), and it leverages a novel reward function designed specifically for AS of disease progression in more than one part of a human body. Our method is end-to-end, relies on multi-modal Deep Learning, and requires no human input at inference time. Throughout an exhaustive experimental evaluation, we show that using RL can provide a higher monetary benefit when compared to state-of-the-art baselines. Finally, our work introduces a novel AS benchmark for RL in medicine, which we make publicly available at https://github.com/Oulu-IMEDS/OACostSensitivityRL.
@@ 1 Introduction
Fig. 1: The workflow of our method is a sequential decision-making process of how frequently each patient should visit the hospital. The state at each time point is associated with the data acquired at the latest patient visit. A reward function is designed to maximize the efficiency of hospital visits by taking into account radiographic changes and hospital visit costs.
@@
To date, there are no effective treatments that can prevent long-term disability and permanent joint damage in KOA [24]. At the latest stages of KOA, a costly and highly invasive intervention, such as total knee replacement (TKR), is delivered to improve the patient’s health and well-being. A TKR surgery may cost up to \$50,000 for each patient [60, 20, 59], and TKR reportedly yields unsatisfactory results in 15-20% of cases [44, 8, 60]. Due to rapid population aging, TKR gradually becomes a financial burden not only for KOA patients [38], but also for the healthcare system [22]. Therefore, KOA in general, requires the next-generation disease-modifying drugs (DMOADs) [52, 65]. Such therapies should aim to slow down the KOA progression, thus, delaying the need for TKR [40, 11].
@@
Despite recent advancements in the field of KOA progression prediction [25, 46, 74, 51, 49], translation of existing models into real-world applications is challenging for several reasons. Firstly, the financial downstream impact of the disease progression models is hard to assess during the model development. As such, the current models do not account for follow-up costs and potential future expenses associated with structural KOA progression when they estimate if a person will develop KOA in the future. Secondly, the existing models [74, 76, 49, 27, 30, 54, 55] largely focus only on the knee-level progression events and overlook KOA’s broader patient-level context, where progression may happen concurrently in both knees. Thirdly, OA severity may increase multiple times during the study time, and there are uncertain future factors that may alter the disease course (e.g. future injuries). Therefore, long-horizon prediction is hard, and data used for such prediction needs to be collected dynamically.
Active Sensing (AS) [84] addresses the question of when to follow up patient and with what tools or modalities an examination has to be conducted. In the case of KOA, one would want for an AS policy to minimize the follow-up costs while maximizing the chances of capturing patient-level progression. Solving such a problem, however, is challenging for a naïve supervised learning (SL) approach, as one needs to consider making predictions to collect data, which would then be used to make new predictions and so forth. This is a hardly differentiable problem of decision-making under uncertainty, which can be solved using e.g. Reinforcement Learning (RL).
In this work, we develop an AS methodology for KOA using RL. We utilize RL to make sequential decisions based on the interactions between a decision-making agent and an environment simulating a clinical trial or another setting where patients need to be observed on an interval basis for a prolonged period of time. Our study makes the following contributions:
@@ 2.1 Problem definition
Our objective is to make optimal, w.r.t. both costs and outcomes, sequential decisions to refer each subject for an examination to capture the disease progression given acquired multi-modal data. We illustrated the decision-making process in Figure 2. Specifically, for each subject, at time point $t\in\{0,\dots,T-1\}$, we predict whether that subject should be examined at the next time point $t+1$. The referral for an examination is considered successful if the disease has progressed between $t$ and $t+1$. We formalize this as the following optimization problem:
$\{a_{0}^{*},\dots,a_{T-1}^{*}\}=\arg\min_{\{a_{1},\dots,a_{T-1}\}}\frac{1}{T-1}\sum_{t=1}^{T-1}J\left(s_{t},a_{t},\{\delta(j)\}_{j=1}^{T};\kappa,\lambda\right),$ (1)
where $a_{t}\in\{0,1\}$ indicates whether to issue a referral for a visit at time-point $t+1$. We assume that the value of the objective $J$ depends on the actions $a_{t}$, states $s_{t}$, and the set of patient-level changes associated with the disease $\{\delta(j)\}_{j=1}^{T}$ (i.e. KOA in the context of this study). We also introduce a hyper-parameter threshold $\kappa$ that defines whether the disease progression is significant (e.g. may correspond to minimal clinically important difference). The combination of changes $\{\delta(j)\}_{j=1}^{T}$ and $\kappa$ defines the disease trajectory. The cost of data acquisition $\lambda$ is treated as a hyper-parameter. We also use $\lambda$ as a basic cost unit multiplier when defining the costs of undetected disease progression.
The optimization of the defined objective in (1) is the problem of sequential decision-making under uncertainty, which we further define through the prism of the Markov Decision Processes (MDP) and RL.
Fig. 2: Data collection as decision-making with active sensing. Every circle represents collected data and the changes in blue color demonstrate the updating process, the darker the later collected time. This figure highlights how present decisions to collect data impact future ones.
@@ 2.2 Reinforcement Learning
In the MDP framework, the sequential decision-making process unfolds as follows [70]. At a time step $t\in\mathbb{N}$, the environment (a “world” in which the agent exists and makes actions) yields a state $s_{t}\in\mathbf{S}$ that is observed by the agent. The agent then responds with an action $a_{t}\in\mathbf{A}$ following the policy $\pi(a_{t}\mid s_{t})$ that assigns a probability to the event associated with selecting action $a_{t}$ at a state $s_{t}$. Subsequently, the environment provides a reward $r_{t}=R(s_{t},a_{t})$ to the agent for taking the action $a_{t}$. The process then repeats until the horizon is reached (i.e. $t=T$). We use the term experiences to refer to a collection of states $s_{t}$, next states $s_{t+1}$, actions $a_{t}$, and rewards $r_{t}$
@@
where $\alpha$ is the learning rate, determining how much newly obtained information replaces the old information. Through interactions with the environment, the policy under Q-learning is supposed to converge after a sufficient number of iterations [79]. The presented update rule is sufficient when the number of states and actions is relatively small. In the context of this study or other real-world cases, the set of states may have an infinite size, making it impossible to store the Q-table. To tackle this issue, one can resort to a functional approximation, such as with a neural network (NN) [72, 47, 48].
@@
In contemporary RL, the Q-function is often approximated by NN, which we denote as $\tilde{Q}(s_{t},a_{t},\theta)$, where $\theta$ represents NN’s parameters. The parameterized Q-function is then commonly called a Q-Network. [47] introduced two networks that allow to follow the Temporal Difference Learning idea [72]: the local Q-Network, approximating $Q(s_{t},\cdot),$ and the target Q-Network, approximating $Q(s_{t+1},\cdot)$. The local Q-Network is optimized by utilizing the mean square error loss to measure the difference between the estimated Q-value and the true one, noted as:
@@
$\pi(a_{t}|s_{t})$ =$=\left\{\begin{matrix}[l]\arg\max_{a}Q(s_{t},a)&\,\ \mathrm{if}\ n_{r}<\varepsilon\\ \textrm{Random action}&\,\ \mathrm{otherwise}\,\end{matrix}\right.$ (6)
@@ 2.3 Proposed method
The progression trajectory of KOA occurs in several stages and simultaneously in both knees. To monitor the disease through sequential actions, we first define the progression events, which in our study, happen multiple times.
In KOA, the decline of JSW measured from plain radiographs is associated with this disease progression. Hence, in this work, we employed the changes in medial fixed JSW@0.250 (fJSW) proposed by [18], which was shown to be a robust and effective biomarker for capturing structural degeneration in KOA [63]. We denote $\delta(t_{r},\zeta)$ as the changes of this biomarker between the reference time point $t_{r}$ and time $\zeta$. When we define progression, the reference point is the closest progression event, starting at the baseline. If $\delta(t_{r},\zeta)$ is greater is greater than a certain threshold, denoted as $\kappa$, we consider $\zeta$ as the time when a progression event occurs and add it to a set of progression events. This set, called $\mathbf{M}_{p}$, includes all significant structural damages observed in either knee. Since a decrease of $0.7\,\mathrm{mm}$ in fJSW is commonly considered the minimum threshold for detecting KOA radiographic progression [19, 23], we used a $\kappa$ of 0.7.
Reward function definition
@@
For each subject, at time point $t$, an agent takes either of the two actions: follow-up or dismiss (i.e. skip examination) at time $t+1$. The rewards for these actions are denoted as $r_{f}$ and $r_{d}$, respectively. There are three possible cases for the correctness of the follow-up action ($a=1$): early visit, timely visit, and late visit. Furthermore, there are two cases for the dismissal action ($a=0$): true dismissal and false dismissal. Mathematically, the reward $r_{t}$ at the time point $t$ corresponding to the follow-up and dismissal actions can be expressed as follows
@@
where $t_{p}$ – the nearest progression time point, $t_{r}$ – reference time point, $\Delta(t_{r},t)=c\delta(t_{r},t)$ – data acquisition utility function, $\lambda\in\mathbb{R}^{+}$ – fixed cost of data acquisition, $\tau:[0,T]\rightarrow\mathbb{R}^{+}$ – sensitivity of the agent’s mistake, $\alpha\in[0,1)$ – parameter to ease the penalty for the late visit, and $\beta<c\kappa$ – positive reward for the correct dismissal action. In our setup, the late dismissal is a false dismissal discounted by $\alpha$ and decreased by $\lambda$. The coefficient $c$ is a scaling factor that converts changes in disease biomarkers to the same measurement units as $\lambda$, and it needs to be devised from the domain knowledge and health economics.
Costs of bilateral structural KOA progression
To set the value of $c$, we rely on the domain knowledge, i.e. (1) the end-stage of the disease is TKR, and (2) if KOA can be detected early, it can presumably be slowed down [82]. From these assumptions, we define $c$ as the monetary loss of $1mm$ of articular cartilage (MLAC), which is calculated as:
$c$ $=\frac{\mathrm{TKR\ cost}\ (\$)}{\mathrm{Mean\ \ JSW\ of\ healthy\ knees}\ (mm)},$ (9)
where the mean JSW can be computed from an observational cohort and TKR cost is pre-defined for a healthcare system where the RL method is deployed.
To quantify the radiographic changes in both knees, we define $\phi_{L}(\zeta)$ and $\phi_{R}(\zeta)$ as the JSW values of the left and right knee, respectively. We then calculate $d_{L}(\zeta)$ and $d_{R}(\zeta)$ that represent the decrease in JSW from the latest visit $t_{r}$ to the assessment time $\zeta$. To mitigate the impact of measurement noise in JSW and handle the rare events of JSW increase, we consider that changes in the JSW w.r.t $t_{r}$ are strictly negative (i.e. JSW is non-increasing):
$d_{L}(t_{r},\zeta)$ $=\max(0,\phi_{L}(t_{r})-\phi_{L}(\zeta)),$ (10)
$d_{R}(t_{r},\zeta)$ $=\max(0,\phi_{R}(t_{r})-\phi_{R}(\zeta)).$ (11)
The function $\delta(\zeta)$, which quantifies the changes at a patient level, is then defined as:
$\delta(t_{r},\zeta)$ =$=\left\{\begin{matrix}d_{LR}(t_{r},\zeta)&\mathrm{if}\ d_{LR}(t_{r},\zeta)\geq\kappa\\ \\ 0&\text{otherwise},\\ \end{matrix}\right.$ (12)
where $d_{LR}(t_{r},\zeta)=\max(d_{L}(t_{r},\zeta),d_{R}(t_{r},\zeta))$. We note that tracking progressions at every knee individually is insufficient since a typical acquisition of X-ray images is bilateral. For that reason, the status of both knees is always updated as newly acquired data. In the unilateral case, the definition of $\delta(\cdot)$ would require minor changes to the environment, and we ablate the effect of single knee modeling in the experimental part of our work.
Weighing early and late follow-up decisions
To incorporate the distance between the time of the follow-up action $t$ and the true OA progression time $t_{p}$ into the reward function, we introduce the modulating function $\tau(t):[0,T]\rightarrow[0,1]$, defined as
@@
We consider two specific cases that require such a modulating function: early follow-up and late follow-up/false dismissal. For the first case, we use an exponential decay function $e^{-\tau(t)}$. Therefore, an increasing $\tau(t)$ reduces the reward that an agent can obtain. For the second case, we penalize late or false dismissal actions, and use the term $-e^{\tau(t)}$. As $\tau(t)$ increases, the negative reward becomes more considerable, discouraging the agent from delaying necessary actions or incorrectly dismissing the patient.

### v2

@@ Abstract
Osteoarthritis (OA) is the most common musculoskeletal disease, which has no cure. Knee OA (KOA) is one of the highest causes of disability worldwide, and it costs billions of United States dollars to the global economy. Prediction of KOA progression has been of high interest to the research community for years, as it can advance treatment development through more efficient clinical trials and improve patient outcomes through more efficient healthcare utilization. Existing approaches for predicting KOA, however, are predominantly static, i.e. consider data from a single time point to predict progression many years into the future, and knee level, i.e. consider progression in a single joint only. Due to these and related reasons, these methods fail to deliver the level of predictive performance, which is sufficient to result in cost savings and better patient outcomes. Collecting extensive data from all patients on a regular basis could address the issue, but it is limited by the high cost at a population level. In this work, we propose to go beyond static prediction models in OA, and bring a novel Active Sensing (AS) approach, designed to dynamically follow up patients with the objective of maximizing the number of informative data acquisitions, while minimizing their total cost over a period of time. Our approach is based on Reinforcement Learning (RL), and it leverages a novel reward function designed specifically for AS of disease progression in more than one part of the human body. Our method relies on multi-modal Deep Learning and requires no human input at test time. Throughout an exhaustive experimental evaluation, we show that using RL can provide a higher monetary benefit when compared to state-of-the-art baselines. Finally, our work introduces a novel AS benchmark for RL in medicine, which we make publicly available at https://github.com/Oulu-IMEDS/OACostSensitivityRL.
@@ 1 Introduction
Fig. 1: The workflow of our active sensing method, is represented as decision-making under uncertainty. The state at each time point is associated with the data acquired at the latest patient visit. A reward function is designed to maximize the efficiency of hospital visits by taking into account radiographic changes and hospital visit costs. The actions set contains two actions – follow-up at time $t$ or skip.
@@
To date, there are no effective treatments that can prevent long-term disability and permanent joint damage in KOA [24]. At the latest stages of KOA, a costly and highly invasive intervention, such as total knee replacement (TKR), is delivered to improve the patient’s health and well-being. A TKR surgery may cost up to \$50,000 for each patient [60, 20, 59], and TKR reportedly yields unsatisfactory results in 15-20% of cases [44, 8, 60]. Due to rapid population aging, TKR gradually becomes a financial burden not only for individuals [38] but also for the healthcare system [22]. Therefore, KOA in general, requires the next-generation disease-modifying drugs (DMOADs) [52, 65]. Such therapies should aim at slowing down the disease progression, thereby delaying the need for TKR [40, 11].
@@
Despite recent advancements in the field of KOA progression prediction [25, 46, 74, 51, 49], translation of existing models into real-world applications is challenging for several reasons. Firstly, the financial downstream impact of the disease progression models is hard to assess during the model development phase. As such, the current models do not account for follow-up costs and potential future expenses associated with structural KOA progression when they estimate if a person will develop KOA in the future. Secondly, the existing models, e.g. [74, 76, 49, 27, 30, 54, 55] largely focus only on the knee-level progression events and overlook KOA’s broader patient-level context, where progression may happen concurrently in both knees. Thirdly, OA severity may increase multiple times during the study time, and there are uncertain future factors that may alter the disease course (e.g. future injuries). Therefore, long-horizon prediction is hard, and one must rethink the paradigm of predicting OA progression.
Active Sensing (AS) [84] addresses the question of when to follow up patient and with what tools or modalities an examination has to be conducted. In the case of KOA, one would want an AS policy to minimize the follow-up costs while maximizing the chances of capturing patient-level progression. Solving such a problem, however, is challenging for a naïve supervised learning (SL) approach, as AS involves making predictions to acquire new data, which in turn informs further predictions, leading to a continuous cycle of adaptive data collection. This is a hardly differentiable problem of decision-making under uncertainty, which can be solved using e.g. Reinforcement Learning (RL) [70].
In this work, we develop an AS methodology for KOA using RL. We utilize RL to make sequential decisions based on the interactions between a decision-making agent and an environment simulating a clinical trial or another setting where patients need to be observed on an interval basis for a prolonged period of time. The overview of our method is depicted in Figure 1, and the specific contributions of our work are:
@@ 2.1 Problem definition
Our objective is to make optimal, w.r.t. both costs and outcomes over a defined time interval $T$, sequential decisions to refer each subject for an examination, to capture the disease progression using the acquired multi-modal data. We illustrated the decision-making process in Figure 2. Specifically, for each subject, at time point $t\in\{0,\dots,T-1\}$, we predict whether that subject should be examined at the next time point $t+1$. The referral for an examination is considered successful if the disease has progressed between $t$ and $t+1$. We formalize this as the following optimization problem:
$\{a_{0}^{*},\dots,a_{T-1}^{*}\}=\operatorname*{arg\,min}_{\{a_{0},\dots,a_{T-1}\}}\frac{1}{T}\sum_{t=0}^{T-1}J\left(s_{t},a_{t},\{y_{i}\}_{i=1}^{T},\lambda\right),$ (1)
where action $a_{t}\in\{0,1\}$ indicates whether to issue a referral for a visit at time-point $t+1$. We assume that the value of the objective $J$ depends on the actions $a_{t}$, states $s_{t}$, and patient-level disease trajectory $\{y_{i}\}_{i=1}^{T}$ (in this study, KOA).The cost of data acquisition $\lambda$ is treated as a hyper-parameter. We also use $\lambda$ as a basic cost unit multiplier when defining the costs of undetected disease progression.
The optimization problem defined in (1) is the problem of learning to make sequential decisions under uncertainty. We further define this problem through the prism of Markov Decision Processes (MDP) and RL.
Fig. 2: Active sensing as decision-making, highlighting how present decisions to collect data impact future ones.
@@ 2.2 Reinforcement Learning
In the MDP framework, the sequential decision-making process unfolds as follows [70]. At a time step $t\in\mathbb{N}$, the environment (a “world” in which the agent exists and makes actions) yields a state $s_{t}\in\mathbf{S}$ that is observed by the agent. The agent then responds with an action $a_{t}\in\mathbf{A}$ following the policy $\pi(a_{t}\mid s_{t})$. The policy assigns a probability to the event associated with selecting action $a_{t}$ at a state $s_{t}$. Subsequently, the environment provides a reward $r_{t}=R(s_{t},a_{t})$ to the agent for taking the action $a_{t}$. The process then repeats until the horizon is reached (i.e. $t=T$). We use the term experiences to refer to a collection of states $s_{t}$, next states $s_{t+1}$, actions $a_{t}$, and rewards $r_{t}$
@@
where $\alpha$ is the learning rate, determining how much newly obtained information replaces the old information. Through interactions with the environment, the policy under Q-learning is supposed to converge after a sufficient number of iterations [79]. The presented update rule is sufficient when the number of states and actions is relatively small. In the context of this study or other real-world cases, the set of states may have an infinite size, making it impossible to store the Q-table.
@@
In contemporary RL, the Q-function is often approximated by a neural network (NN) [72, 47, 48], which we denote as $\tilde{Q}(s_{t},a_{t},\theta)$, where $\theta$ represents NN’s parameters. The parameterized Q-function is then commonly called a Q-network. [47] introduced a method with two networks that allow to follow the Temporal Difference Learning idea [72]: the local Q-Network, approximating $Q(s_{t},\cdot),$ and the target Q-Network, approximating $Q(s_{t+1},\cdot)$. The local Q-Network is optimized by utilizing the mean square error loss to measure the difference between the estimated Q-value and the true one, noted as:
@@
$\pi(a_{t}\mid s_{t})$ =$=\left\{\begin{matrix}[l]\arg\max_{a}Q(s_{t},a)&\,\ \mathrm{if}\ n_{r}<\varepsilon\\ \textrm{Random action}&\,\ \mathrm{otherwise}\,\end{matrix}\right.$ (6)
@@ 2.3 Proposed method
In this section, we introduce a procedure to construct a ground-truth set of progression events for individual subjects using the decline in medial fixed JSW@0.250111Hereinafter, the two terms JSW and JSW@0.25 are used interchangeably., a robust and effective biomarker for capturing structural degeneration in KOA [63]. We use $\phi_{L}(\zeta)$ and $\phi_{R}(\zeta)$ to respectively denote the measured JSWs of the left and the right knees at every observational time point $\zeta\in\{0,\dots,T\}$. The changes in left and right JSWs at time point $\zeta$ relative to a reference time point $\zeta_{r}$, where $\zeta_{r}<\zeta$, are denoted by $d_{L}(\zeta_{r},\zeta)$ and $d_{R}(\zeta_{r},\zeta)$, respectively. The following procedure aims to generate the set of progression events $\mathbf{M}_{p}$, which was done once before training the RL policy.
Initially, we set the first reference time point to the baseline examination, denoted by $\zeta_{r}=0$, and let $\mathbf{M}_{p}$ be empty. At each time step $\zeta\in\{1,\dots,T\}$, we calculate $d_{L}(\zeta_{r},\zeta)$ and $d_{R}(\zeta_{r},\zeta)$ as follows:
$d_{L}(\zeta_{r},\zeta)$ $=\max(0,\phi_{L}(\zeta_{r})-\phi_{L}(\zeta)),$ (7)
$d_{R}(\zeta_{r},\zeta)$ $=\max(0,\phi_{R}(\zeta_{r})-\phi_{R}(\zeta)).$ (8)
Next, we define the patient-level KOA worsening at $\zeta$ relative to $\zeta_{r}$ as:
$d_{LR}(\zeta_{r},\zeta)=\max(d_{L}(\zeta_{r},\zeta),d_{R}(\zeta_{r},\zeta)).$ (9)
If $d_{LR}(\zeta_{r},\zeta)$ exceeds a threshold $\kappa$, a KOA progression event is recorded, and we add $\zeta$ to $\mathbf{M}_{p}$. We then set $\zeta$ as the new reference time point (i.e., $\zeta_{r}:=\zeta$) and continue the procedure at $\zeta+1$. If the threshold $\kappa$ is not exceeded, the procedure also advances to $\zeta+1$. Ultimately, when $\zeta$ reaches $T$, we terminate the process and obtain the progression set $\mathbf{M}_{p}$ containing the time points of all the progression events.
Monetization of bilateral structural KOA progression
In this section, we aim to harmonize two semantically distinct units – bilateral structural changes (measured in mm) and acquisition costs (measured in \$currency-dollar\\$\$) – in our reward function. Leveraging domain knowledge, we propose a novel approach to convert JSW reduction to a monetary value. Specifically, we assume that (1) the end-stage of the disease is TKR, and (2) if KOA is detected early, it can presumably be slowed down [82]. From these assumptions, we define $c$ as the monetary loss of $1\mathrm{mm}$ of articular cartilage (MLAC), calculated as:
$c$ $=\frac{\mathrm{TKR\ cost}\ (\$)}{\mathrm{Mean\ \ JSW\ of\ healthy\ knees}\ (\mathrm{mm})},$ (10)
where the mean JSW can be computed from an observational cohort and TKR cost is pre-defined for a healthcare system where the RL method is deployed. To this end, we can map bilateral structural changes (see eq. (9)) into financial costs as
$\delta(\cdot,\cdot)=cd_{LR}(\cdot,\cdot).$ (11)
Reward function
Given the set of progression events $\mathbf{M}_{p}$ and the monetized radiographic changes, we here develop a novel reward function to optimize the AS objective stated in Section 2.1. This reward function is designed to provide our agent with monetary costs of its decisions by combining the costs from bilateral structural changes in the knees and acquisition costs.
@@
For each subject, at time point $t$, an agent takes one of two actions – to follow up or to dismiss (i.e. skip examination) at time $t+1$. The rewards for these actions are denoted as $r_{f}$ and $r_{d}$, respectively. The reference point $t_{r}$ is updated to the latest visit, ensuring that the reward calculations are based on the most recent state of the subject’s condition. By comparing $t+1$ with $t_{p}$ – the nearest progression time point $t_{p}$ extracted from the patient-level progressions set $\mathbf{M}_{p}$, we derived five possible scenarios. Namely, there are three cases for the correctness of the follow-up action ($a=1$): early visit, timely visit, and late visit and two cases for the one of the dismissal action ($a=0$): true dismissal and false dismissal.
Mathematically, the reward $r_{t}$ at the time point $t$ corresponding to the follow-up and dismissal actions can be expressed as follows:
@@
where $\Delta(t_{r},\cdot)$ is the data acquisition utility function, representing the cost of significant changes relative to $t_{r}$, formulated as:
$\Delta(t_{r},\cdot)$ =$=\left\{\begin{matrix}\delta(t_{r},\cdot)&\mathrm{if}\ \delta(t_{r},\cdot)\geq c\kappa\\ \\ 0&\text{otherwise},\\ \end{matrix}\right.$ (14)
$\lambda\in\mathbb{R}^{+}$ is the fixed cost of data acquisition. The reward for a late visit is equivalent to that for a false dismissal, yet discounted by $\alpha\in[0,1)$ and decreased by $\lambda$. In addition, $\beta<c\kappa$ is the positive reward for the correct dismissal action.
To incorporate the difference between the time of making decision $t$ and the true KOA progression time $t_{p}$ into the reward function, we introduce the modulating function $\tau:[0,T]\rightarrow[0,1]$, defined as:
@@
As such, we consider two specific cases that require such a modulating function: early follow-up and late follow-up/false dismissal. For the first case, we use an exponential decay function $e^{-\tau(t)}$. Therefore, an increasing $\tau(t)$ reduces the reward that an agent can obtain. For the second case, we penalize late or false dismissal actions and use the term $-e^{\tau(t)}$. As $\tau(t)$ increases, the negative reward becomes more considerable, discouraging the agent from delaying necessary actions or incorrectly dismissing the patient.